9. (5 marks) Consider a Gamma random variable, Y ~ Ganzma(α = n/2, β). Find the moment- generating function of U = c Y. If U ~ , what is c?
Answer #1

9)

Here Ysim Ga(alpha=rac{n}{2},eta)

So, MGF of Y = M_{Y}(t)=(1-eta t)^{-rac{n}{2}}

Now, U=cY, so MGF of U is M_{U}(t)=E(e^{tU})=E[e^{t[cY]}]=M_{Y}(ct)=(1-eta ct)^{-rac{n}{2}} --------(1)

Also, we know the if Usim chi _{n}^{2} , then MU (t) = (1-20-2 ----------(2)

On comparing (1) and (2), we get, eta c=2Rightarrow c=rac{1}{eta}